OpenStudy (anonymous):

1. If (a,b) is a solution of the system of equations {P=0 and Q=0}, show that it must also be a solution of the system of equations {P=0 and Q+kP=0}. 2. If (a,b) is a solution of the system of equations {P=0 and Q+kP=0}, show that it must also be a solution of the system of equations {P=0 and Q=0}.

OpenStudy (anonymous):

Where are you stuck?

OpenStudy (anonymous):

1) if (a,b) is a solution of (P=0 and Q =0) , that is P(a,b) =0 and Q (a,b) =0 since P(a,b) =0, k*P(a,b) =0 =k*0=0

OpenStudy (anonymous):

I don't even know where to begin. I am completely lost.

OpenStudy (anonymous):

on the other hand (Q + kP =0) means (Q+kP) (a,b) =0 . this is Q+ kP "apply on (a,b), it is NOT (Q+kP)* (a,b)

OpenStudy (anonymous):

You must understand this concept to step up

OpenStudy (anonymous):

Thank you, but I just wanted to know what would the final answer be?

OpenStudy (anonymous):

For example: P(x) = x^2-1 , then if x =1 is a solution of P(x), then P(1) =0

OpenStudy (anonymous):

YOur problem is a "Show" problem, it doesn't have final answer.

OpenStudy (anonymous):

You have to do some logic to get the credit. In other words, you have to PROVE it

OpenStudy (anonymous):

I give you the whole proof for the first one, you do the second one, OK?

OpenStudy (anonymous):

Again, thank you.

OpenStudy (anonymous):

(a,b) is a solution of P=0 and Q =0 (given) . That is P(a,b) =0 , hence k*P(a,b) =k*P(a,b) =0 hence \(Q(a,b) =0 ~~(given)\\k*P(a,b) =0~~(above)\\-------------------\\Q(a,b)+kP(a,b) =0+0=0\\(Q+kP)(a,b)=0\) that shows (a,b) is a solution of \((Q+kP)\) Dat sit